The owner of a firm expects to make a profit of $12,000 for each of the next two years and to be able to sell the firm at the end of the second year for $30,000. The owner of the firm believes that the appropriate discount rate for the firm is 14 percent. Calculate the value of the firm.
Answer #1

value of the firm means the present value of the cash flows.

Present value = A(P/A, i, n) + F(P/F, i, n)

                      = 12,000(P/A, 14%, 2) + 30,000(P/F, 14%, 2)

                      = 12,000(1.647) + 30,000(0.7695)

                      = 19,764 + 23,085

                      = $42,849

Thus, the value of the firm is $42,849